Find a two digit number such that the tens digit exceeds the unit digit by 3 and the number is 4 less than the sum of the squares of its digits.
Answer by mananth(16946)

let units digit be y
ten's digit be x
...
x=y+3
x-y=3
...
10x+y= x^2+y^2-4
x^2+y^2-y-10x-4 =0
plug value of x
(y+3)^2+y^2-y-10(y+3)-4=0
y^2+6y+9+y^2-y-10y-30-4=0
2y^2-5y-25=0
2y^2-10y+5y-25=0
2y(y-5)+5(y-5)=0
(2y+5)(y-5)=0
y=5
x=8
the number is 85